A foreign exchange trader based in the US, authorized to borrow $450,000 or its foreign currency equivalent faces the following quotes:

Spot rate: $1.3000/pound

Six Month Forward: $1.3085/pound

US Interest Rate: 3.0% per annum

UK Interest Rate: 2.0% per annum

Is covered Interest arbitrage possible, and if so, how much profit can the trader make via 1 covered interest arbitrage transaction? Please show all steps and work.
Answer #1
1) Forward premium on the pound = 1.3085/1.3000-1 = 0.65%
Difference in 6 months rate = 1.5%-1.0% = 0.50%
2) As the forward premium and the difference in
interest rates are not equal, there is scope for
covered interest rate arbitrage.
Since the forward premium is more, it would be
advantageous to borrow in the currency having
higher interest rate and to invest the proceeds in
the currency having lower interest rate.
The steps involved would be:
a) Borrow $450000 for 6 months; the loan will have
a maturity of 450000*1.015 = $          4,56,750
b) Convert the $450000 into GBP at spot to get
450000/1.3 = £          3,46,154
c) Invest the GBP for 6 months to get 346154*1.01 = £          3,49,615
after 6 months
d) Sell forward GBP349615.38 at 1.3085 to get 349615*1.3085 = $          4,57,471
3) After 6 months:
Close the GBP deposit and convert it realize $457471;
then pay the loan amount with interest amounting
to $456750, thereby netting a riskless profit of
457471-456750 = $                   721